OpenStudy (anonymous):

Solve the inequality: 3x + 2 ≤ 17 Answer: X>=-5

OpenStudy (anonymous):

@mathmath333

OpenStudy (texaschic101):

3x + 2 <= 17 --- subtract 2 from both sides 3x <= 17 - 2 3x <= 15 -- divide both sides by 3 x >= 5 (not -5)

OpenStudy (mathmath333):

\(\Large 3x\leq 15\\ \Large x\leq \dfrac{15}{3}\\ \Large x\leq 5\) the inequality sign only chandes when u multiply or divide both sides by \(\Large -1\)